OpenStudy (anonymous):

I need help!! Proof that d/dx e^x=e^x..but I must use limit h-->infinity (1+1/h)^h

OpenStudy (anonymous):

How do you want to prove..it can easily be proved through the first property of derivative...f(x+h)-f(x)/h exists,where h ->0

OpenStudy (anonymous):

I know but the professor says that we must use the property of the limit that defines e.

OpenStudy (anonymous):

The key steps are: 1) We must use the second definition of e = lim (h->0)(1 + h)^(1/h). 2) The definition of e implies as h->0, e ~ (1 + h)^(1/h). Raise both sides to the h power to obtain e^h ~ (1 + h).

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Im confused how e^h=(1+h), the graphs are totally different

OpenStudy (anonymous):

We're considering the case as h -> 0. Of course e^h and 1 + h are very different as h -> infinity. However, as h -> 0, e^h and 1 + h are closer and close. At h = 0, e^h = 1+ h. Since h <> 0, we can only evaluate as h -> 0, or e^h ~ 1 + h. But this is sufficient to evaluate the limit without error. Looking at the graph, you will see they are closer and close as h -> 0.

OpenStudy (anonymous):

that is suffice for a mathematical argument?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Professor doesn't accept it lol

OpenStudy (anonymous):

i think this depends entirely on your definition of \(e^x\)

OpenStudy (anonymous):

one definition is that \(e\) is the number such that \(\lim_{x\to 0}\frac{e^x-1}{x}=1\) if that is your definition then you get it right away.. if not i guess you have to do something else

OpenStudy (anonymous):

this problem is from three days ago. what is it doing here?

OpenStudy (anonymous):

Sorry i still need help. (The professor is saying that the limit def. of e (limit (1+ 1/h)^h) needs to be used to imply d/dx e^x=e^x

OpenStudy (experimentx):

\[ \lim_{h \rightarrow \infty, \Delta x \rightarrow 0 } \frac{ \left ( 1 + \frac{x+\Delta x}{h} \right )^h - \left ( 1 + \frac{x}{h} \right )^h }{ \Delta x}\]

OpenStudy (experimentx):

Expand this binomially \[ \left ( 1 + \frac x h + \frac{\Delta x} h\right )^h = \left ( 1 + \frac{x}{h}\right )^h + h \left ( 1 + \frac{x}{h}\right )^{h-1} \frac{\Delta x}{h} \\+ \text{ Terms containing higher power of }\Delta x^2\] Subtracting you get, \[ \lim_{h, \Delta x--}\frac{\left ( 1 + \frac{x}{h}\right )^{h-1} \Delta x + O(n > 1)} {\Delta x} \\ = \lim_{h \rightarrow \infty }\left ( 1 + \frac{x}{h}\right )^{h-1} = e^x\]

OpenStudy (anonymous):

He told me to google finding limits by substitution then he said " e^h-1=U "

OpenStudy (unklerhaukus):

\[e=\sum\limits_{n=0}^\infty\frac{1}{n!}\] \[e^x=\sum\limits_{n=0}^\infty\left(\frac{1}{n!}\right)^x \]

OpenStudy (experimentx):

\[ e^x=\sum\limits_{n=0}^\infty\left(\frac{x^n}{n!}\right) \]

OpenStudy (anonymous):

Where are you guys using limit by substitution?

OpenStudy (experimentx):

well ... my answer was from binomial expansion.

OpenStudy (unklerhaukus):

how come you have an \(n\) as an exponent to \(x\) in your equation and i dont @experimentX ?

OpenStudy (experimentx):

e^x = 1 + x/1 + x^2/2! + x^3/3! + ... so on x^n/n!

OpenStudy (anonymous):

I dont see how d/dx e^x=e^x is implied

OpenStudy (unklerhaukus):

expand the sum @jk_16

OpenStudy (unklerhaukus):

each term is the derivative of the term next to in in the sequence

OpenStudy (experimentx):

well this works pretty fine. lim Δx -> 0 e^(x + Δx) - e^x/Δx = e^x

OpenStudy (anonymous):

he wants limit by substitution

OpenStudy (experimentx):

i can't see where you can use substitution.